Richard

Question 2


In a study of hypertension and optimal treatment conducted by the National Heart Institute, 10,000 patients had a mean systolic blood pressure (BP), 𝝁 = πŸπŸ”πŸ mm Hg and standard deviation, 𝝈 = πŸπŸ“ mm Hg. Assume the systolic blood pressure is normally distributed.



d. If 60 random samples each of size 30 are drawn from this population, determine:


i. the sampling distribution of the mean systolic blood pressure.


ii. the probability that the of mean systolic blood pressure between 140 and 165 mm Hg.


Hi, regarding question d ( ii) based from#180930 / how you calculate the answer P(0.4478) and also the answer 0.5376, can show me the exact calculations. tq


When x = 165


\frac{x-\mu}{\sigma}=\frac{x-161}{42.4264}=0.4478 \\ \therefore P(Z\le 0.4478) = 0.5376

Οƒ

xβˆ’ΞΌ

​

=

42.4264

xβˆ’161

​

=0.4478

∴P(Z≀0.4478)=0.5376





1
Expert's answer
2021-04-25T16:01:14-0400

We have,

"\mu = 161"

"\sigma = 25"

Since, the systolic blood pressure is normally distributed

i) "f(x) = \dfrac{1}{\sigma \sqrt{2\pi}}e^-\dfrac{1}{2}(\dfrac{x- \mu}{\sigma})^2"


"f(x) = \dfrac{1}{25\sqrt{2\pi}}e^-\dfrac{1}{2}(\dfrac{x- 161}{25})^2"


ii) "P(140<X<165)"


"= P(\dfrac{140-165}{25}<Z< \dfrac{165-161}{25})"


"=P(-1<Z<0.16)"

"= P(-1<Z<0)+P(0<Z<0.16)\\= P(0<Z<1)+P(0<Z<0.16)"

"= 0.0398+0.636"

"= 0.67"